OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality −x − 3 < 13 + 3x. −x − 3 < 13 + 3x −4x < 16 x > −4 A)−4x − 3 > 13 B)2x − 3 < 18 C) −4x − 3 < 13 D) 2x − 3 > 18

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks!!! i got more... can you help? @aajugdar

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

C put -2 on the other side you will get x<=-8

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

you r welcome ^_~

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7 − 9x − (x + 12) <= 25? x −3 x −3 −3 x A number line with a closed circle on negative 3 and shading to the right.

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

@Nurali not a way she said?

OpenStudy (anonymous):

i already did C: want me to make these into more questions and add you to them so i can choose yours as best responses so it will benefit you? (:

OpenStudy (anonymous):

wait a bit your options they were wrong rephrase them please

OpenStudy (anonymous):

x −3 x −3 −3 x these

OpenStudy (anonymous):

x >= −3 x <= −3 −3 <= x A number line with a closed circle on negative 3 and shading to the right.

OpenStudy (anonymous):

then the answer is A because its incorrect

OpenStudy (anonymous):

is C -3>=x or -3<=x

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

okay then any more?

OpenStudy (anonymous):

i made a new question